# Covariant Fractons and Weitzenböck Torsion

### v1

@@ 3 Review of teleparallel gravity
$T_{\mu}{}^{\varrho}{}_{\nu}(\Gamma)$ $=\Gamma_{\mu}{}^{\varrho}{}_{\nu}-\Gamma_{\nu}{}^{\varrho}{}_{\mu}=0,$ (3.1b)
$R_{\mu\nu}{}^{\varrho}{}_{\sigma}(\Gamma)$ $=\partial_{\mu}\,\Gamma_{\nu}{}^{\varrho}{}_{\sigma}-\partial_{\nu}\,\Gamma_{\mu}{}^{\varrho}{}_{\sigma}+\Gamma_{\mu}{}^{\varrho}{}_{\tau}\,\Gamma_{\nu}{}^{\tau}{}_{\sigma}-\Gamma_{\nu}{}^{\varrho}{}_{\tau}\,\Gamma_{\mu}{}^{\tau}{}_{\sigma}.$ (3.1c)
@@
There is an alternative but equivalent formulation of General Relativity (at least classically and without boundary terms), known as teleparallel gravity, in which one chooses a connection $W_{\mu}{}^{\varrho}{}_{\nu}$, compatible with the metric, such that the dynamical variables are encoded in the torsion and the curvature vanishes Einstein1928 ; Weitzenbock1928 :
@@
$T_{\mu}{}^{\varrho}{}_{\nu}(W)$ $=W_{\mu}{}^{\varrho}{}_{\nu}-W_{\nu}{}^{\varrho}{}_{\mu},$ (3.3b)
$R_{\mu\nu}{}^{\varrho}{}_{\sigma}(W)$ $=\partial_{\mu}\,W_{\nu}{}^{\varrho}{}_{\sigma}-\partial_{\nu}\,W_{\mu}{}^{\varrho}{}_{\sigma}+W_{\mu}{}^{\varrho}{}_{\tau}\,W_{\nu}{}^{\tau}{}_{\sigma}-W_{\nu}{}^{\varrho}{}_{\tau}\,W_{\mu}{}^{\tau}{}_{\sigma}=0.$ (3.3c)
While in the usual formulation of General Relativity the gravitational interaction is due to the curvature of spacetime, in teleparallel gravity the gravitational interaction is mediated by a Lorentz-like force in a flat spacetime, as in Yang-Mills theories deAndrade:1997gka . Indeed, while in the former case the action is linear in the curvature, in the latter one the action is quadratic in the torsion.
Since the spacetime manifold is supposed to be flat, then it is possible to choose a system of locally inertial coordinates valid everywhere. In other words, the frame or inverse vielbein $e^{\mu}{}_{a}$ is a set of $d$ vectors in $d$ dimensions in the tangent space of the manifold such that, when it is evaluated in a point, it forms a basis for the tangent space in that point, for all the points of the manifold. In such a case the manifold is said parallelisable, and it is possible to choose a connection such that the vielbein is parallel or covariantly constant
$\partial_{\mu}\,e_{\nu}{}^{a}-W_{\mu}{}^{\varrho}{}_{\nu}\,e_{\varrho}{}^{a}=0,$ (3.4)
@@
which consistently has vanishing curvature, since it is the Cartan-Maurer form out of the vielbein. [^5] The spin connection corresponding to the Weitzenböck connection is zero (Weitzenböck condition). See A for more details. Moreover, the condition (3.4) implies the compatibility condition (3.3a), since
@@
$\Gamma_{\mu}{}^{\varrho}{}_{\nu}=W_{\mu}{}^{\varrho}{}_{\nu}+K_{\mu}{}^{\varrho}{}_{\nu}=W_{\mu}{}^{\varrho}{}_{\nu}+\tfrac{1}{2}\,(T_{\mu\nu}{}^{\varrho}+T_{\nu\mu}{}^{\varrho}-T_{\mu}{}^{\varrho}{}_{\nu}),$ (3.11)
where the tensor $-K_{\mu}{}^{\varrho}{}_{\nu}$ is the contorsion. Using the vanishing of the curvature of the Weitzenböck connection and the relation (3.11), one can write the Riemann tensor in terms of the contorsion in the following way:
@@
$e\,R=-\partial_{\mu}\,(2\,e\,T^{\mu\nu}{}_{\nu})-e\,(\tfrac{1}{4}\,T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}+\tfrac{1}{2}\,T^{\mu\nu\varrho}\,T_{\mu\varrho\nu}-T_{\mu\nu}{}^{\nu}\,T^{\mu\varrho}{}_{\varrho}),$ (3.15)
@@
where we set $\kappa=8\,\pi\,G=1$.
@@ 4 Review of Møller-Hayashi-Shirafuji theory
The relation (3.16) shows that the Einstein-Hilbert action consists in a linear combination of three quadratic scalar contractions of the Weitzenböck torsion with precise relative coefficients. In Moller1961a ; Moller1961b ; Moller1978 ; Hayashi:1967se ; Hayashi:1979qx the theory defined by the same action with arbitrary coefficients is considered:
$S_{\textsc{mhs}}=\frac{1}{2}\int\mathrm{d}^{d}x\,e\,(\alpha_{1}\,T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}+\alpha_{2}\,T^{\mu\nu\varrho}\,T_{\mu\varrho\nu}+\alpha_{3}\,T^{\mu\nu}{}_{\nu}\,T_{\mu\varrho}{}^{\varrho}).$ (4.1)
@@
is captured by formally requiring the local Lorentz invariance, which indeed is the necessary requirement in the vielbein formulation of General Relativity, beside diffeomorphism invariance, in order to recover the right number of degrees of freedom Cho:1975dh .
The mhs action is the most general quadratic action in the Weitzenböck torsion, since there are only three possible independent quadratic scalar contractions of the torsion. Indeed, the torsion has three irreducible components:
$T_{\mu\varrho\nu}\rightarrow(1,1)\otimes(1)=(1,1,1)\otimes(2,1)_{t}\otimes(1),$ (4.3)
where $(\lambda_{1},\lambda_{2},\dots,\lambda_{p})$ denotes a Young tableau with $p$ rows and $\lambda_{i}$ boxes in the $i$-th row, $i=1,\dots,p$; and $t$ denotes the traceless part. $(1,1,1)$ is the totally antisymmetric part $T_{[\mu\nu\varrho]}$, and $(1)$ is the unique trace $T_{\mu\nu}{}^{\nu}$. Namely, the contraction out of two traces is $T_{\mu\nu}{}^{\nu}\,T^{\mu\varrho}{}_{\varrho}$; the contraction out of two totally antisymmetric tensors is $T^{[\mu\nu\varrho]}\,T_{[\mu\nu\varrho]}$, which is proportional to $T^{\mu\nu\varrho}\,T_{[\mu\nu\varrho]}$; the scalar contraction involving the remaining irreducible component can be replaced by $T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}$; and the second contraction can be replaced by $T^{\mu\nu\varrho}\,T_{\mu\varrho\nu}$, since
@@
$-\tfrac{1}{2}\,g_{\mu\nu}\,(\alpha_{1}\,T^{\alpha\beta\gamma}\,T_{\alpha\beta\gamma}+\alpha_{2}\,T^{\alpha\beta\gamma}\,T_{\alpha\gamma\beta}-\alpha_{3}\,T^{\alpha\beta}{}_{\beta}\,T_{\alpha\gamma}{}^{\gamma}-2\,\alpha_{3}\,D_{\alpha}\,T^{\alpha\beta}{}_{\beta}),$ (4.7a)
$E_{\mu\nu}^{\text{(anti)}}=$ $\,(\tfrac{\alpha_{2}}{2}-\alpha_{1})\,(T_{[\mu\nu]}{}^{\alpha}\,T_{\alpha\beta}{}^{\beta}+D_{\alpha}\,T_{[\mu\nu]}{}^{\alpha})-(\alpha_{2}+\tfrac{\alpha_{3}}{2})\,(T_{\mu\alpha\nu}\,T^{\alpha\beta}{}_{\beta}+D_{\alpha}\,T_{\mu}{}^{\alpha}{}_{\nu}).$ (4.7b)
@@
where [^7] The inverse relations read: $f_{\mu\nu\varrho}=\tfrac{1}{3}\,(T_{\mu\nu\varrho}+T_{\varrho\mu\nu}-2\,T_{\nu\varrho\mu})$ and $t_{\mu\nu\varrho}=\tfrac{1}{3}\,(T_{\mu\nu\varrho}+T_{\varrho\mu\nu}+T_{\nu\varrho\mu})$.
@@ 5 Linearisation of Møller-Hayashi-Shirafuji theory
$\frac{1}{2}\int\mathrm{d}^{d}\,x\,(\alpha_{1}\,\mathring{T}^{\mu\nu\varrho}\,\mathring{T}_{\mu\nu\varrho}+\alpha_{2}\,\mathring{T}^{\mu\nu\varrho}\,\mathring{T}_{\mu\varrho\nu}+\alpha_{3}\,\mathring{T}^{\mu\nu}{}_{\nu}\,\mathring{T}_{\mu\varrho}{}^{\varrho}).$ (5.13)
@@
So, one has to study the local $s$-cohomology of ghost number zero on the jet space of $\overline{h}_{\mu\nu}$ and its derivatives. This is done in Appendix B. The result is that the $s$-cohomology is generated by the derivatives of the antisymmetrised combination $\partial_{[\mu}\,\overline{h}_{\varrho]\nu}$, which is the linearised torsion $\mathring{T}_{\mu\nu\varrho}$. Thus, one has to consider the most general action quadratic in $\mathring{T}_{\mu\nu\varrho}$, which is a combination of the three independent quadratic scalar contractions $\mathring{T}^{\mu\nu\varrho}\,\mathring{T}_{\mu\nu\varrho}$, $\mathring{T}^{\mu\nu\varrho}\,\mathring{T}_{\mu\varrho\nu}$, $\mathring{T}^{\mu\nu}{}_{\nu}\,\mathring{T}_{\mu\varrho}{}^{\varrho}$ with arbitrary coefficients, and this is the action (5.13).
@@
where $\mathring{\Omega}_{\mu\nu}=-\mathring{\Omega}_{\nu\mu}$ is anticommuting. Redefining $\tilde{\Omega}_{\nu\mu}=:\partial_{\mu}\,\mathring{\xi}_{\nu}-\partial_{\nu}\,\mathring{\xi}_{\mu}+2\,\mathring{\Omega}_{\nu\mu}$, one realises that the transformation of the antisymmetric part becomes $s_{\textsc{fp}}\,b_{\mu\nu}=\tilde{\Omega}_{\mu\nu}$, $s_{\textsc{fp}}\,\tilde{\Omega}_{\mu\nu}=0$, so that $(b_{\mu\nu},\tilde{\Omega}_{\nu\mu})$ is a trivial doublet, and the local $s_{\textsc{fp}}$-cohomology does not depend on it Brandt:1989rd ; Piguet:1995er . On the other hand, the transformation of the symmetric part is the same as before $s_{\textsc{fp}}\,h_{\mu\nu}=s\,h_{\mu\nu}$. Therefore, the local $s_{\textsc{fp}}$-cohomology on the jet space of $\{h_{\mu\nu},\mathring{\xi}_{\mu},b_{\mu\nu},\tilde{\Omega}_{\mu\nu}\}$ and their derivatives is equivalent to the local $s$-cohomology on the jet space of the starting field space $\{h_{\mu\nu},\mathring{\xi}_{\mu}\}$.
@@
6 Embedding covariant fractons in linearised MHS theory
In this section we show that all the solutions of the equations of motion of covariant fracton gauge theory (2.5) or (2.9) are solutions of the equations of motion of linearised mhs theory (5.5)–(5.6).
The space of the solutions of covariant fracton gauge theory modulo the gauge invariance (fracton moduli space) is the submanifold $\mathscr{M}_{\text{fr}}$ in the space of fields identified by the equations of motion (2.5) or (2.9),
$\mathscr{M}_{\text{fr}}=\{f^{\text{(fr)}}_{\mu\nu\varrho}(x)\,:\,E^{\text{(fr)}}_{\mu\nu}(f^{\text{(fr)}})=0\}$ (6.1)

### v3

@@ 3 Review of teleparallel gravity
[…]
@@
There is an alternative but equivalent formulation of General Relativity (at least classically and without boundary terms), known as teleparallel equivalent of General Relativity, in which one chooses a connection $W_{\mu}{}^{\varrho}{}_{\nu}$, compatible with the metric, such that the dynamical variables are encoded in the torsion and the curvature vanishes Einstein1928 ; Weitzenbock1928 :
@@
[…]
While in the usual formulation of General Relativity the gravitational interaction is due to the curvature of spacetime, in teleparallel formulation the gravitational interaction is mediated by a Lorentz-like force in a flat spacetime, as in Yang-Mills theories deAndrade:1997gka . Indeed, while in the former case the action is linear in the curvature, in the latter one the action is quadratic in the torsion.
Denoting with $e_{\mu}{}_{a}$ the components of the vielbein and with $e^{\mu}{}_{a}$ the components of the frame or inverse vielbein, it is possible to choose a connection $W_{\mu}{}^{\varrho}{}_{\nu}$ such that the vielbein is parallel or covariantly constant, that is the following equation holds:
$\partial_{\mu}\,e_{\nu}{}^{a}-W_{\mu}{}^{\varrho}{}_{\nu}\,e_{\varrho}{}^{a}=0.$ (3.4)
@@
which consistently has vanishing curvature, since it is the Cartan-Maurer form out of the vielbein. [^5] The spin connection corresponding to the Weitzenböck connection is zero (Weitzenböck condition). See Appendix A for more details. Moreover, the condition (3.4) implies the compatibility condition (3.3a), since
@@
[…]
where the tensor $K_{\mu}{}^{\varrho}{}_{\nu}$ is the contorsion. Using the vanishing of the curvature of the Weitzenböck connection and the relation (3.11), one can write the Riemann tensor in terms of the contorsion in the following way:
@@
[…]
@@
where we set $\kappa=8\,\pi\,G=1$. [^6] This proves the dynamical equivalence of the usual formulation of General Relativity and of the teleparallel equivalent, since it is not necessary to assume the existence of a globally defined frame Cederwall:2021xqi . Although not manifest in teleparallel formulation, the precise combination in (3.16) is locally Lorentz invariant (see next Sections and Cederwall:2021xqi ), and the inverse frame serves as a gauge field. Thus, the global extension of the frame is ensured by suitably choosing the gauge redundancy in the overlaps of the coordinates patches.
@@ 4 Review of Møller-Hayashi-Shirafuji theory
The relation (3.16) shows that the Einstein-Hilbert action consists in a linear combination of three quadratic scalar contractions of the Weitzenböck torsion with precise relative coefficients. In Moller1961a ; Moller1961b ; Moller1978 ; Hayashi:1967se ; Hayashi:1979qx the theory defined by the same action with arbitrary coefficients is considered,
[…]
@@
is captured by formally requiring the local Lorentz invariance, which indeed is the necessary requirement in the vielbein formulation of General Relativity, beside diffeomorphism invariance, in order to recover the right number of degrees of freedom Cho:1975dh . [^8] Since in mhs theory local Lorentz-invariance is spoiled, unless (4.2) are met, one has to assume the parallelisability of spacetime a priori, that is the existence of a frame which serves as basis for the tangent bundle, which trivialises Hayashi:1979qx .
The mhs action is the most general quadratic action in the Weitzenböck torsion, since there are only three possible independent quadratic scalar contractions of the torsion. Indeed, the torsion has three irreducible components: [^9] See note 2 for the notation of Young tableaux.
$T_{\mu\varrho\nu}\rightarrow(1,1)\otimes(1)=(1,1,1)\otimes(2,1)_{t}\otimes(1).$ (4.3)
$(1,1,1)$ is the totally antisymmetric part $T_{[\mu\nu\varrho]}$, and $(1)$ is the unique trace $T_{\mu\nu}{}^{\nu}$. Namely, the contraction out of two traces is $T_{\mu\nu}{}^{\nu}\,T^{\mu\varrho}{}_{\varrho}$; the contraction out of two totally antisymmetric tensors is $T^{[\mu\nu\varrho]}\,T_{[\mu\nu\varrho]}$, which is proportional to $T^{\mu\nu\varrho}\,T_{[\mu\nu\varrho]}$; the scalar contraction involving the remaining irreducible component can be replaced by $T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}$; and the second contraction can be replaced by $T^{\mu\nu\varrho}\,T_{\mu\varrho\nu}$, since
@@
[…]
@@
where [^10] The inverse relations read $f_{\mu\nu\varrho}=\tfrac{1}{3}\,(T_{\mu\nu\varrho}+T_{\varrho\mu\nu}-2\,T_{\nu\varrho\mu})$ and $t_{\mu\nu\varrho}=\tfrac{1}{3}\,(T_{\mu\nu\varrho}+T_{\varrho\mu\nu}+T_{\nu\varrho\mu})$.
@@ 5 Linearisation of Møller-Hayashi-Shirafuji theory
[…]
@@
So, one has to study the local $s$-cohomology of ghost number zero on the space of $\overline{h}_{\mu\nu}$ and its derivatives, treated as independent variables (the so-called jet space). This is done in Appendix B. The result is that the $s$-cohomology is generated by the derivatives of the antisymmetrised combination $\partial_{[\mu}\,\overline{h}_{\varrho]\nu}$, which is the linearised torsion $\mathring{T}_{\mu\nu\varrho}$. Thus, one has to consider the most general action quadratic in $\mathring{T}_{\mu\nu\varrho}$, which is a combination of the three independent quadratic scalar contractions $\mathring{T}^{\mu\nu\varrho}\,\mathring{T}_{\mu\nu\varrho}$, $\mathring{T}^{\mu\nu\varrho}\,\mathring{T}_{\mu\varrho\nu}$, $\mathring{T}^{\mu\nu}{}_{\nu}\,\mathring{T}_{\mu\varrho}{}^{\varrho}$ with arbitrary coefficients, and this is the action (5.13).
@@
where $\mathring{\Omega}_{\mu\nu}=-\mathring{\Omega}_{\nu\mu}$ is anticommuting. Redefining $\tilde{\Omega}_{\nu\mu}=:\partial_{\mu}\,\mathring{\xi}_{\nu}-\partial_{\nu}\,\mathring{\xi}_{\mu}+2\,\mathring{\Omega}_{\nu\mu}$, one realises that the transformation of the antisymmetric part becomes $s_{\textsc{fp}}\,b_{\mu\nu}=\tilde{\Omega}_{\mu\nu}$, $s_{\textsc{fp}}\,\tilde{\Omega}_{\mu\nu}=0$, so that $(b_{\mu\nu},\tilde{\Omega}_{\nu\mu})$ is a trivial doublet, and the local $s_{\textsc{fp}}$-cohomology does not depend on it Brandt:1989rd ; Piguet:1995er . On the other hand, the transformation of the symmetric part is the same as before $s_{\textsc{fp}}\,h_{\mu\nu}=s\,h_{\mu\nu}$. Therefore, the local $s_{\textsc{fp}}$-cohomology on the space of $\{h_{\mu\nu},\mathring{\xi}_{\mu},b_{\mu\nu},\tilde{\Omega}_{\mu\nu}\}$ and their derivatives is equivalent to the local $s$-cohomology on the space of the starting field space $\{h_{\mu\nu},\mathring{\xi}_{\mu}\}$ and derivatives.
@@
6 Particle spectrum in covariant fracton gauge theory
The space of the solutions of covariant fracton gauge theory $\mathscr{M}_{\text{fr}}$ (moduli space) is the submanifold in the space of fields $\{h^{\text{(fr)}}_{\mu\nu}\}$, identified by the equations (2.5) or (2.10), modulo the gauge invariance, [^12] The author thanks C. Imbimbo, who suggested how to expand this and the next section which, as stated by the brst rule in (2.1), [^13] In this section $\lambda$ is a simple gauge parameter. We maintain the same notation of the corresponding ghost because this should not cause confusion. amounts to linearised longitudinal diffeomorphisms,
$\delta_{\text{fracton}}\,h_{\mu\nu}^{\text{(fr)}}=\partial_{\mu}\,\partial_{\nu}\,\lambda,\;\;\forall\;\alpha,\beta.$ (6.1)